OpenStudy (anonymous):

Help!?! Simplify!?!? y > -3x − 4.

OpenStudy (anonymous):

B. 13/14

OpenStudy (anonymous):

first do (2/3 * 6/7) = 12/21 then do (5/8 divided by 1 and 3/4).. we know 1 and 3/4 = 7/4 so to divide, you multiply by the reciprocal. so (5/8 / 7/4) becomes (5/8 * 4/7) = 20/56 last step is to add (12/21) + (20/56) = 13/14